The Fresh Connections makes a special run of Orange/Mango juice with extra Vitamin C each week of the summer for sports camps. They complete this run as a special weekend shift producing 6906 cases PET bottles of the special juice to meet their demand for the camps. They need workers to come in to load the filled PET bottles into the special cases used for the sports camps. They run a special shift of 6 hours. On average, a worker can package 119 cases per hour worked. Your policy is that all workers will work a full shift with the exception of one worker that will just make up the difference to get to the cases needed for the week, i.e., you can have a fractional person if they aren't needed to work a full shift. Your workers have indicated that they believe they should get paid for a full shift even if they aren't needed for the full shift. If you pay the workers $11.94 per hour, how much more will it cost for an average week if you agree to pay all of the workers for the full shift?

In the proposed scenario you will need to round up the number of workers to a full worker. Don't round anything else until you reach the end of the problem and then round to two (2) decimal places.

Hint: The answer is 23.4787
Answer #1

Total cases to be packaged = 6906 cases

Number of cases that can be packed per hour by a worker = 119 cases

Number of hours required to pack 6906 cases= 6906/119 = 58.03361345 hours

Number of hours in a shift =6 hours

Number of workers required in a shift = 58.03361345/6 = 9.67227

Hence 10 workers are required in a shift to pack 6906 cases in which 9 workers would work 6 hours and one worker would work 4.03361345 hours to complete 58.03361345 hours.

Labor charge per hour= $11.94

If the employer pays full amount for all the workers, total payment = 11.94*10*6= $716.4

If the employer pay as per number of hours worked, the payment=(9*11.94*6)+(1*11.94*4.03361345)

=644.76+48.161345 = $692.921345

The extra amount that will cost to the employer = 716.4-692.921344 = 23.478655

=$23.4787